# Trial and Trust: Addressing Byzantine Attacks with Comprehensive Defense Strategy

### v1

@@ 3.1  First method: BANT
In this section, we introduce our method designed to counter Byzantine attacks – Algorithm 1, termed Byzantines ANTidote (BANT).
Our method relies on the core idea of assigning trust scores to devices, a technique gaining popularity in defending against attacks (Cao et al., 2020; Yan et al., 2024). We integrate this with the concept of a trial function by aggregating the stochastic gradients $g_{i}^{t}$ of devices with their respective weights $w_{i}^{t}$. To find the latter, we firstly calculate the contribution coefficients for each worker $i$ at each step: ${}\theta_{i}^{t}=\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma g_{i}^{t})~{}$. These coefficients show how the $i$-th device affects the convergence. If $\theta_{i}^{t}>0$, the stochastic gradient minimizes the trial function, and is assigned some weight. Otherwise, it is assigned zero weight (Line 10 in Algorithm 1). We ensure non-negativity with $\left[\theta_{i}^{t}\right]_{0}=\max\{\theta_{i}^{t},0\}$ and normalize to provide a total weight of 1. To address stochastic gradient instability, which can increase the loss function, we introduce a momentum parameter for the weights (Line 10). If all gradients increase the loss, they are given zero weights, stopping the minimization process even without Byzantine devices in the network. By adding momentum, we achieve a more stable convergence in practice. This allows previous good gradients to influence current weights, even if a device receives a small or zero weight in the current iteration. An indicator in the step (Line 14) ensures that gradients maximizing the trial function are ignored, guaranteeing its minimization at each step.
@@
2:  Parameters: Stepsize $\gamma>0$, momentum $\beta\in[0,1]$
@@
11:     if each $[\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma g_{i}^{t})]_{0}=0$ then
@@
16:  Output: $\frac{1}{T}\sum\limits_{t=0}^{T-1}x^{t}$
Now we show the convergence results.
@@
Under Assumptions 2.1, 2.2(b), 2.3, 2.4 with $\delta_{2}\leqslant\frac{1}{12}$, 2.5, for solving the problem (1), after $T$ iterations of Algorithm 1 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}\cdot\frac{4n}{\beta G}+6L\gamma\sigma^{2}+3\delta_{1}+4\zeta(N).$
The first two terms in the result of Theorem 3.1 replicate the findings from the standard SGD analysis, up to constant factors. The last term, which depends on $\zeta(N)$, is of special interest. The function $\zeta(N)$ reflects the relationship between $f_{1}$ and the trial loss $\hat{f}$ and represents an approximation error. The dependence of $\zeta(N)$ on $N$ is natural: the larger $N$ is, the smaller the error. More precisely, for our function $f$, this error is $\zeta(N)=\mathcal{\widetilde{O}}\left(\frac{1}{N}\right)$ (see Lemma C.1). Although this error degrades convergence, it is actually very common for machine learning tasks. In particular, the original learning problem like (1) is often replaced by its Monte Carlo approximation (Johnson & Zhang, 2013; Defazio et al., 2014; Allen-Zhu, 2018), and the resulting problem is often referred to as the empirical risk minimization (Shalev-Shwartz & Ben-David, 2014). This replacement also leads to an error, but is often not mentioned, considering it a classical and common phenomenon. Finally, $\delta_{1}$ is a typical term which represents data similarity (Assumption 2.4) and is unavoidable in the presence of Byzantines (Wu et al., 2020; Karimireddy et al., 2021a; Gorbunov et al., 2022).
Since our approach bears resemblance with the Zeno method (Xie et al., 2019) in its use of the trial functions, we must mention that we found some issues in their proofs. Specifically, in Theorem 1 of (Xie et al., 2019), the authors incorrectly apply the expectation operator when deriving their recursion. They sample a trial function from the full dataset and use $\mathbb{E}[\hat{f}^{t}(x)]=f(x)$, which is true for a random point. However, in the case of the point $x^{t+1}$, they make an error. Since they sample $\hat{f}^{t}$ in every iteration, the point $x^{t+1}$ depends on the sample $\hat{f}^{t}$, leading to $\mathbb{E}[\hat{f}^{t}(x^{t+1})\mid x^{t}]\neq f(x^{t+1})$. If one attempts to carry through the conditional expectation without considering the full expectation, it becomes impossible to enter the recursion and achieve convergence with respect to the function $f$ itself. To resolve this, we explicitly bound the gradient discrepancy $|\nabla f_{1}(x^{t})-\nabla\hat{f}(x^{t})|$, leveraging the server’s trial function sampling to ensure statistical convergence as sample size grows. This difference in Theorem 3.1 is represented by the discussed $\zeta(N)$. Now we present the final convergence.
@@
$\frac{1}{T}\sum\limits_{t=1}^{T-1}$ $\mathbb{E}\|\nabla f(x^{t})\|^{2}=\mathcal{O}\left(\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}{\beta GT}+\frac{\sigma\cdot\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]\cdot Ln}}{\sqrt{\beta GT}}+\delta_{1}+\zeta(N)\right).$
If we consider the first two terms in the convergence estimates from Corollary 3.2, the only difference from the classical SGD convergence results Moulines & Bach (2011); Stich (2019) is the additional factor $\frac{n}{G}$, but the rate is asymptotically optimal. The proof of this corollary, the main theorem, as well as the result for the strongly-convex objective, are presented in Appendix, D.
3.2  Second method: AutoBANT
Despite all advantages of the BANT algorithm, it has some imperfections connected to the mechanism of assigning trust scores. The parameter $\beta$ has a negative effect. While it helps honest clients maintain trust scores despite occasional bad gradients, it also allows Byzantine devices to retain their weights during attacks. To combat this, we add an indicator for the trial function reduction (the indication of the device being Byzantine at the considered iteration). However, this limits the theoretical applicability of the method to non-convex problems, prevalent in modern machine learning. To resolve these limitations, we present our second method, called AUxiliary Trial Optimization for Byzantines ANTidote (AutoBANT), formalized as Algorithm 2.
@@ 3.2  Second method: AutoBANT
10:      $\omega^{t}\approx\arg\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)$
11:      $x^{t+1}=x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$
@@
13:  Output: $\frac{1}{T}\sum\limits_{t=0}^{T-1}x^{t}$
The idea of assigning weights to devices as a part of the optimization process has recently gained popularity in federated learning. For instance, in many works, it leads to improved solution quality (Li et al., 2023; Tupitsa et al., 2024), or is used in more specific settings such as personalized learning (Mishchenko et al., 2023). We propose to adapt this to Byzantine optimization by optimizing the functionality of $\hat{f}$ with respect to weights calculated after each algorithmic step (Line 10). To solve the minimization problem, we can use various methods, e.g., Mirror Descent (Beck & Teboulle, 2003; Allen-Zhu & Orecchia, 2014):
$\omega^{k+1}~{}=\arg\underset{\omega\in\Delta_{1}^{n}}{\min}$$\Bigl{\{}\eta\left\langle\nabla_{\omega}\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{k}g_{i}^{t}\right),\omega\right\rangle+\mathcal{KL}(\omega\|\omega^{k})\Bigr{\}},$
where $\mathcal{KL}(\cdot\|\cdot)$ denotes the Kullback-Leibler divergence. The error of solving this is bounded by $\delta$:
$\vspace{-2mm}\left|\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)-\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}\right)\right|\leqslant\delta.$
After solving this auxiliary problem, we produce an actual model update using the optimized weights (Line 11). In light of the proposed method, the question of the cost of implementing such an optimal scheme comes to the forefront. Note that the computational complexity of solving this subproblem at each iteration is only $\mathcal{O}(\nicefrac{{\log n}}{{\delta^{2}}})$ (Beck & Teboulle, 2003), which is not critical.
@@
Under Assumptions 2.1, 2.2(c), 2.3, 2.4 with $\delta_{2}<\frac{1}{12}$, 2.5, for solving the problem 1, after $T$ iterations of Algorithm 2 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}$$\leqslant$ $\frac{4\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}+3\delta_{1}+\frac{6L\gamma}{G}\sigma^{2}+2\zeta(N)+\frac{4\delta}{\gamma}.$
@@
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f(x^{t})\right\|^{2}=\mathcal{O}\left(\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]L}{T}+\delta_{1}+\zeta(N)+\frac{\sigma\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]L}}{\sqrt{TG}}+\left(L+\frac{\sqrt{TL}\sigma}{\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]G}}\right)\delta\right).$
Detailed proofs are presented in Appendix E. Let us discuss the obtained result. In the first and second terms, we see that the method converges the same way as the standard SGD only on honest workers Ghadimi & Lan (2013); Ghadimi et al. (2016). It turns out that we just throw out all Byzantines and this result is almost optimal and unimprovable. As in the case of Algorithm 1, a term responsible for the approximation error $\zeta(N)$ appears in the convergence. By comparing with the result of Corollary 3.2, we improve the rate through a more advanced aggregation mechanism, specifically by removing the factor $\frac{n}{\beta G}$ from the main term and achieving a reduction in variance by a factor of $G$. However, we incurred an additional error $\delta$, which can be seen as a trade-off for solving the subproblem. Furthermore, our approach applies to a broader class of non-convex functions. Addressing the dependence of the stepsize on the number of Byzantine workers, this choice is based on theoretical analysis of the worst-case scenario, considering the number of Byzantine workers. If this number is unknown, setting the minimum possible value of 1 eliminates this dependency.
@@ 4  Extensions
Byzantine robust optimization, as discussed earlier, lacks a solid theoretical foundation in several formulations that are the most applicable in the real world. In this work, we address this gap. This section provides a brief overview of the scenarios to which we extend our analysis.
@@ 4.1  Local methods
The main idea is that each device performs a predefined number of local steps. Then the aggregation of gradients and mutual updates of model parameters, initialized by the server, take place. This reduces the number of communication rounds. However, it affects the convergence proportionally to the length of the communication round. Complete updates are performed only at specific iterations: $t=t_{k\cdot l}\text{ for some }k=\overline{0,\left\lfloor\nicefrac{{T}}{{l}}\right\rfloor}$. During the remaining iterations, we simply perform local updates of the points using the rule $x_{i}^{t+1}=x_{i}^{t}-\gamma g_{i}^{t}$. This approach ensures that communication overhead is minimized while maintaining efficient convergence (see Appendix G).
@@ 4.2  Partial participation
It occurs when only a subset of clients actively participates in the training process during each communication round (Yang et al., 2021), allowing clients to join or leave the system. This approach is beneficial in scenarios like mobile edge computing. However, it poses challenges such as incomplete model updates and potential degradation in model performance due to missed contributions from inactive clients (Wang & Ji, 2022; Li et al., 2022). Our methods adapt to the scenario of partial participation due to the assignment of trust scores to devices explicitly participating in training at the considered iteration. Furthermore, it is crucial to account for the minimum number of nodes participating in training across all iterations. Specifically, we analyze $\widetilde{G}(t)=\min_{t\leqslant T}G(t)$, where $G(t)$ denotes the set of active honest workers at iteration $t$. The full version is in Appendix H.

### v5

@@ 4.1 First method: Bant
In this section, we introduce our method, termed Byzantine ANTidote (Bant) – Algorithm 1. Our method relies on the core idea of assigning trust scores to devices. We integrate this with the concept of a trial function by aggregating the stochastic gradients $g_{i}^{t}$ of devices with their respective weights $w_{i}^{t}$. To find the latter, we first calculate the contribution coefficients for each worker $i$ at each step: $\theta_{i}^{t}=\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma g_{i}^{t})~$. These coefficients demonstrate how the $i$-th device affects convergence. If $\theta_{i}^{t}>0$, the stochastic gradient minimizes trial loss and is assigned a weight. Otherwise, it is assigned a weight of zero (Line 10 in Algorithm 1). We ensure non-negativity with $\left[\theta_{i}^{t}\right]_{0}=\max\{\theta_{i}^{t},0\}$ and normalize to provide a total weight of 1.
To address stochastic gradient instability, which can increase the loss function, we introduce a momentum parameter for the weights (Line 10). If all gradients increase the loss, they are assigned zero weights, thereby stopping the minimization process even in the absence of Byzantine devices in the network.
@@
2:Parameters: Stepsize $\gamma>0$, momentum $\beta\in(0,1]$
@@
11:  if each $\hat{f}(x^{t})-\hat{f}(x^{t}-\gamma g_{i}^{t})]_{0}=0$ then
@@
16:Output: $\frac{1}{T}\sum\nolimits_{t=0}^{T-1}x^{t}$
By adding momentum, we achieve a more stable convergence in practice. This allows previous favorable gradients to influence current weights, even if a device receives a small or zero weight in the current iteration. An indicator in the step (Line 14) ensures that gradients maximizing trial loss are ignored, thereby guaranteeing minimization at each step. We also define $G={\min}_{t}G(t)$ as the minimum, taken over all iterations, of the number of honest workers at each iteration. Now we are ready to proceed with the theoretical results.
@@
Under Assumptions 3.1, 3.2(b), 3.3, 3.4 with $\delta_{2}\leqslant\frac{1}{12}$, 3.5, for solving the problem (1), after $T$ iterations of Algorithm 1 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\textstyle{\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}}\textstyle{\leqslant}$ $\textstyle{\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}\cdot\frac{4n}{\beta G}+6L\gamma\sigma^{2}}$
$\textstyle{+3\delta_{1}+4\zeta(N).}$
The first two terms in the result of Theorem 4.1 replicate the findings from the standard SGD analysis, up to constant factors. The last term, which depends on $\zeta(N)$, is of special interest. The function $\zeta(N)$ reflects the relationship between $f_{1}$ and the trial loss $\hat{f}$ and represents an approximation error. The dependence of $\zeta(N)$ on $N$ is natural: the larger $N$, the smaller the error becomes. More precisely, for our function $f$, this error is $\zeta(N)=\mathcal{\widetilde{O}}\left(\frac{1}{N}\right)$ (see Lemma C.1 in Appendix). Although this error degrades convergence, it is common in machine learning tasks. In particular, the original learning problem, like (1), is often replaced by its Monte Carlo approximation (NIPS2013_ac1dd209; defazio2014saga; allen2018katyusha), and the resulting problem is commonly referred to as empirical risk minimization (shalev2014understanding). This replacement also leads to an error. Finally, $\delta_{1}$ is a typical term that represents data similarity (Assumption 3.4) and is unavoidable in the presence of Byzantines (wu2020federated; karimireddy2021byzantine; gorbunov2022secure).
Since our approach resembles Zeno (xie2019zeno) in its use of the trial function, we should mention that we found some issues in their proofs. In Theorem 1 of (xie2019zeno), the authors incorrectly apply the expectation operator when deriving their recursion. They sample a trial function from the full dataset and use $\mathbb{E}[\hat{f}^{t}(x)]=f(x)$, which is valid for a random point. However, in the case of point $x^{t+1}$, a mistake was made. Since they sample $\hat{f}^{t}$ in every iteration, the point $x^{t+1}$ depends on the sample $\hat{f}^{t}$, leading to $\mathbb{E}[\hat{f}^{t}(x^{t+1})\mid x^{t}]\neq f(x^{t+1})$. By carrying the conditional expectation without considering the full expectation, it becomes impossible to enter the recursion and achieve convergence regarding the function $f$ itself. In turn, we explicitly bound the gradient discrepancy $|\nabla f_{1}(x^{t})-\nabla\hat{f}(x^{t})|$, leveraging trial function sampling to ensure convergence as the sample size increases. This difference in Theorem 4.1 is represented by the discussed $\zeta(N)$.
@@
$\textstyle{\frac{1}{T}\sum\limits_{t=1}^{T-1}}$ $\textstyle{\mathbb{E}\|\nabla f(x^{t})\|^{2}=\mathcal{O}\biggl(\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]Ln}{\beta GT}}$
$\textstyle{+\frac{\sigma\cdot\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]\cdot Ln}}{\sqrt{\beta GT}}+\delta_{1}+\zeta(N)\biggr).}$
If we consider the first two terms in the convergence estimates from Corollary 4.2, the only difference from the classical SGD convergence results (NIPS2011_40008b9a; stich2019unified) is the additional factor $\frac{n}{G}$, however, the rate is asymptotically optimal. The proofs and results for the strongly-convex objective can be found in Appendix D.
4.2 Improving theoretical estimates: AutoBant
Despite practical advantages of Bant, it has some theoretical imperfections related to the mechanism of assigning trust scores. While parameter $\beta$ helps honest clients maintain trust scores despite occasional bad gradients, it also enables Byzantine devices to retain their weights during attacks. To combat this, we add an indicator for the trial function reduction (the indication of the device being Byzantine at the considered iteration). However, this limits the theoretical applicability of the method to non-convex problems prevalent in modern machine learning. To resolve these limitations, we present our second method, called AUxiliary Trial Optimization for Byzantines ANTidote (AutoBant), Algorithm 2.
The idea of assigning weights to devices as part of the optimization process has gained popularity in federated learning. For instance, in many works, it leads to improved solution quality (li2023revisiting; tupitsa2024federated), or it is used in more specific settings, such as personalized learning (mishchenko2023partially). We propose adapting this to Byzantine-robust learning by optimizing $\hat{f}$ regarding weights calculated after each algorithmic step (Line 10).
@@ 4.2 Improving theoretical estimates: AutoBant
10:   $\omega^{t}\approx\arg\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma\sum\nolimits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)$
11:   $x^{t+1}=x^{t}-\gamma\sum\nolimits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}$
@@
13:Output: $\frac{1}{T}\sum\nolimits_{t=0}^{T-1}x^{t}$
To solve the minimization problem, we can use various methods, e.g., Mirror Descent (beck2003mirror; allen2014linear):
$\textstyle{\omega^{k+1}~=\arg\underset{\omega\in\Delta_{1}^{n}}{\min}}$ $\textstyle{\Bigl\{\eta\left\langle\nabla_{\omega}\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{k}g_{i}^{t}\right),\omega\right\rangle}$
$\textstyle{+\mathcal{KL}(\omega\|\omega^{k})\Bigr\}},$
where $\mathcal{KL}(\cdot\|\cdot)$ denotes the Kullback-Leibler divergence. The error in solving this is bounded by $\delta$:
$\textstyle{\left|\underset{\omega\in\Delta_{1}^{n}}{\min}\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}g_{i}^{t}\right)-\hat{f}\left(x^{t}-\gamma\sum\limits_{i=1}^{n}\omega_{i}^{t}g_{i}^{t}\right)\right|\leqslant\delta.}$
After solving this auxiliary problem, we produce an actual model update using the optimized weights (Line 11). In light of the proposed method, the question of the cost of implementing such an optimal scheme comes to the forefront. Note that the computational complexity of solving this subproblem at each iteration is only $\mathcal{O}(\nicefrac{{\log n}}{{\delta^{2}}})$ (beck2003mirror), which is not critical. For the following theoretical analysis, we assume the minimization subproblem can be solved to arbitrary precision, neglecting errors in our estimates. In practice, however, the convergence rate is directly influenced by the accuracy of this solution. A numerical study of this effect is presented in Table 6 in Appendix. Now we are ready to present the main theoretical result of this section.
@@
Under Assumptions 3.1, 3.2(c), 3.3, 3.4 with $\delta_{2}<\frac{1}{12}$, 3.5, for solving the problem 1, after $T$ iterations of Algorithm 2 with $\gamma\leqslant\frac{1}{13L}$, the following holds:
$\textstyle{\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\|\nabla f(x^{t})\|^{2}\leqslant}$ $\textstyle{\frac{4\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]}{\gamma T}+3\delta_{1}}$
$\textstyle{+\frac{6L\gamma}{G}\sigma^{2}+2\zeta(N).}$
@@
$\textstyle{\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}\left\|\nabla f(x^{t})\right\|^{2}=\mathcal{O}\Biggl(}$ $\textstyle{\frac{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]L}{T}+\delta_{1}+\zeta(N)}$
$\hskip 5.69054pt\textstyle{+~\frac{\sigma\sqrt{\mathbb{E}\left[\hat{f}(x^{0})-\hat{f}(\hat{x}^{*})\right]L}}{\sqrt{TG}}\Biggr).}$
Detailed proofs are presented in Appendix E. In the first and second terms, we observe that the method converges in the same pace as the standard SGD only with honest workers (ghadimi2013stochastic; ghadimi2016mini). It turns out that we throw out all Byzantines, and this result is nearly optimal and unimprovable. As in Algorithm 1, a term responsible for the approximation error $\zeta(N)$ appears. Compared with the result of Corollary 4.2, we improve the rate through a more advanced aggregation mechanism. We remove the factor $\frac{n}{\beta G}$ from the main term and achieve a decrease in variance by a factor of $G$. However, an additional error $\delta$ is incurred, which can be viewed as a trade-off for solving the subproblem. Furthermore, our approach applies to a broader class of non-convex functions. Addressing the dependence of the stepsize on the number of Byzantines, this choice is based on the theoretical analysis of the worst-case scenario, considering the number of Byzantines. If this number is unknown, setting the minimum possible value of 1 eliminates this dependency.
@@ 5 Extensions
Byzantine robust optimization, as discussed earlier, lacks a solid theoretical foundation in several real-world settings. We address this gap. This section provides a brief overview of the scenarios to which we extend our analysis.
@@ 5.1 Local methods
The main idea is that each device performs a predefined number of local steps. Then the aggregation of gradients and the mutual updates of model parameters, initialized by the server, take place. This reduces the number of communication rounds. However, it affects the convergence proportionally to the length of the communication round. Complete updates are performed only at specific iterations: $t=t_{k\cdot l}\text{ for some }k=\overline{0,\left\lfloor\nicefrac{{T}}{{l}}\right\rfloor}$. During the remaining iterations, we perform local updates using the rule $x_{i}^{t+1}=x_{i}^{t}-\gamma g_{i}^{t}$. This approach ensures that communication overhead is minimized while maintaining efficient convergence.
@@ 5.2 Partial participation
It occurs when only a subset of clients actively participates in the training process during each communication round (yang2021achieving), allowing clients to join or leave the system. This approach is beneficial in scenarios such as mobile edge computing. However, it poses challenges such as incomplete model updates and potential degradation in model performance due to missed contributions from inactive clients (wang2022unified; li2022federated). Our methods adapt to the scenario of partial participation by assigning trust scores to devices explicitly participating in training during the considered iteration. Furthermore, it is crucial to account for the minimum number of nodes participating in training across all iterations. Specifically, we analyze $\widetilde{G}(t)=\min_{t\leqslant T}G(t)$, where $G(t)$ denotes the set of active honest workers at iteration $t$.